OpenStudy (anonymous):

In bernoulli's THEOREM! large area has high pressure and low speed,where as the formula of pressure is P=F/A?

OpenStudy (turingtest):

Bernoulli's principle is a direct result of conservation of energy, and the formula for pressure you posted above. Summing up the energy in the system we have\[E=K+U_g+W=\frac12mv^2+mgh+Fd\]subbing mass=density*volume, and distance=(volume)/(cross-sectional area):\[E=\frac12\rho Vv^2+\rho Vgh+F(\frac VA)=\frac12\rho Vv^2+\rho Vgh+(\frac FA)V\]\[=\frac12\rho Vv^2+\rho Vgh+pV\]divide by rho V, and note that energy per unit volume is a constant for an incompressible fluid.\[\frac{v^2}2+gh+\frac p\rho=\text{constant}\]